# Leveraging LLM Tutoring Systems for Non-Native English Speakers in Introductory CS Courses

### v1

@@ 1. Introduction
In computing education, being a non-native English speaker (NNES) can present learning challenges, especially for those outside anglophone countries or with limited English proficiency, due to the predominance of English-based materials (Becker, 2019). The most popular programming languages (e.g., Java, Python, C) use English for their keywords and API naming conventions (Hanselman, 2008; Chistyakov, 2017; Igawa et al., [n.d.]). English computer terminology used in classroom settings is one of the main factors negatively affecting NNES student performance in programming classes (Keen and Etzkorn, 2009), predicting approximately 50% of final exam performance when controlling for English proficiency (Alaofi and Russell, 2022). Moreover, English proficiency is strongly correlated with programming performance (Qian and Lehman, 2016; Rauchas et al., 2006; Wibowo and Tandiono, 2022; Pudyastuti et al., 2014), further widening the gap for NNES students. In addition to programming languages being primarily based in English, most online resources for learning programming (e.g., documentation, online Q&A forums, tutorials) are also predominantly in English. These skewed linguistic accommodations make a certain level of English proficiency increasingly necessary to learn computer programming (Hanselman, 2008). This de facto prerequisite creates a significant hurdle for NNES individuals who may have limited English proficiency (Guo, 2018). While tailoring the delivery of materials for NNES students has been suggested (Becker et al., 2022), the practical implementation of this strategy has traditionally been a challenge. However, large language models (LLMs) may now provide an opportunity to better serving NNES computing students as the models having been trained on 100+ natural languages (Xue et al., 2020). Indeed, with the multilingual capabilities of LLMs and their past implementation in computing classrooms (Denny et al., 2024c; Prather et al., 2023; Liu and M’Hiri, 2024; Liu et al., 2024; Woodrow et al., 2024), this technology might help alleviate the de facto primarily English setting by allowing NNES students to easily obtain assistance in their native language.
@@
With the rapid rise of LLMs, computing educators face the challenge of integrating this technology into the classroom (Becker et al., 2023). The study by Denny et al. suggests that LLMs can enhance the accessibility and scalability of high-quality educational content while reducing educators’ workloads (Denny et al., 2023a). One implementation of this technology is through tutoring systems that provide guidance, rather than direct answers to programming questions, helping users to reach solutions independently (Kazemitabaar et al., 2024; Liffiton et al., 2023).
@@
LLM tutoring systems have shown positive effects for students in the classroom (Kazemitabaar et al., 2024; Liffiton et al., 2023; Zapata et al., 2023). Coupled with this, state-of-the-art LLMs that have been trained on 100+ languages (Xue et al., 2020) can perform extremely well when translating between languages (Jordan et al., 2024). Due to this, we hypothesize that NNES students may especially benefit from the use of LLM tutoring systems to ask for help—including in their native language. To understand this, we asked the following research questions:
  1. (1)
@@
2. Related Work
2.1. Non-Native English Speaker Stressors and LLM Assistance
@@
Past research has highlighted several challenges faced by NNES students in academic environments, particularly in fields like computer science (Keen and Etzkorn, 2009; Hanselman, 2008; Guo, 2018; Alaofi and Russell, 2023). Alaofi and Russell highlighted various NNES student difficulties in CS1, including reading, understanding, or using error messages, understanding or using hints and tips from Integrated Development Environments in English, and general English language difficulties (Alaofi and Russell, 2023). Due to a mixture of these difficulties, NNES students often experience linguistic anxiety (Pappamihiel, 2001) and communication apprehension (Sobotka and Raman, 2020) which can be exacerbated in an academic setting where technical terminology is prevalent. These hurdles, coupled with the embarrassment that NNES students often feel when asking for help (Agarwal et al., 2022), can impact their learning process and willingness to seek necessary assistance. Additionally, tutors have reported additional time required to support NNES students (Hope and Chamberlain, 2003), which is in line with tutors reporting overall difficulty in implementing best tutor practices due to time constraints (Riese et al., 2021).
2.2. LLMs in the Computing Classroom
@@
The increasing use of LLMs in computing classrooms (Denny et al., 2024c; Prather et al., 2023; Liu and M’Hiri, 2024; Liu et al., 2024; Denny et al., 2024d; Woodrow et al., 2024; Denny et al., 2024a) presents a valuable resource for students and new opportunities to support learning. State of the art models provide an alternative avenue for help seeking and can therefore aid learning while potentially reducing the workload on tutors (Denny et al., 2024b; Liu et al., 2024), although potential challenges to learning also exist (Prather et al., 2024). Recent work also points to particular benefits for NNES students (au2 et al., 2024). Within computing education, NNES students have been found to grasp computing concepts more quickly when using their native language (Dasgupta and Hill, 2017). Given the strong performance of LLMs in non-English languages (Whitehouse et al., 2023; Jiao et al., 2023; Holtermann et al., 2024a) and the fact that many models are trained on over 100 languages (Xue et al., 2020), LLMs could be particularly beneficial for students studying in their preferred language. LLMs have the potential to effectively bridge language gaps in computing education, which is especially important given that learning to interact productively with LLMs is likely to become an essential skill for all students (Denny et al., 2023b).
@@
This leads to our study in which we use the CodeHelp LLM tutor (Liffiton et al., 2023; Sheese et al., 2024). CodeHelp is an open-source LLM-powered tool designed with guardrails to offer on-demand assistance to computing students without directly providing code solutions. We chose this tool as prior research indicated that it was well-received by students and instructors alike, and presented users with a simple interface (Liffiton et al., 2023). CodeHelp was praised for its availability, help with resolving programming errors, easy deployment, and its supplemental support (Liffiton et al., 2023).
3. Method
3.1. LLM Tutor Deployment and Study Context
@@
In this study, we use the CodeHelp system—an open source LLM-powered tutor that includes built-in guardrails to produce general guidance toward a solution to a programming problem while being prevented from providing direct solutions (Liffiton et al., 2023). Students interact with CodeHelp in a simple form, providing the programming language of interest (e.g., Java), an extract of the code related to the doubt, the error message, and the student’s question. The form can be seen in Figure 1. The written question is what will be analyzed in this study.
Figure 1. CodeHelp user interface illustrating input fields
@@
The implementation of CodeHelp aimed to enhance students’ learning experiences by providing an accessible, AI-driven resource for supplementary instruction and problem-solving support. CodeHelp leverages the OpenAI API with a choice to use the GPT-3.5-Turbo or the GPT-4o models. In our study, we used the GPT-3.5-Turbo model for financial reasons. CodeHelp was introduced to the course via a class forum post from the instructor. After some use with the system, we noticed that the CodeHelp system would always provide help in English, regardless of the language of the query. Due to the nature of this study, we coordinated with the CodeHelp developer to update the system to reply with the same natural language in which a question was asked, a characteristic known as language fidelity (Holtermann et al., 2024b).
@@
Participants for this study were recruited from the pool of 321 students enrolled in the aforementioned accelerated course. All students in the course had access to the CodeHelp tool, but were not required to participate in our study surveys. CodeHelp was available to students for a total of five weeks. Each student independently signed up to the CodeHelp system, ensuring individualized tracking of interactions and usage patterns. Enrollment in the study was voluntary, and students were informed of the study’s purpose, procedures, and potential benefits prior to participating. In order to incentivize survey completion, the professor of the course offered extra credit points for completing surveys but students could still chose to participate in the study or not. In total, 80 students participated in the first survey and 162 students participated in the second survey, for a total of 170 unique students. Of these 170 students, 109 were native English speakers and 61 were non-native English speakers (as seen on Table 1). All participation procedures were in accordance with our approved human subjects protocol.
3.2. Survey Design and Administration
@@ 3.3. Data Analysis
To answer RQ1 around interactions with the LLM tutor, the enrollment and usage rates were analyzed using a $\chi^{2}$ test to explore differences between NNES and NES students. Specifically, we compared the proportions of NNES and NES students who enrolled in CodeHelp and those who actively used the platform post-enrollment. Additionally, we calculated the unique users by hour of a given day to analyze usage patterns for each group. For natural language use, the $langdetect$ Python package was used to identify the languages asked. A manual inspection of all non-English queries was also carried out to verify the correct language identification. We examined the frequency and provide descriptive statistics on the languages used to ask queries to CodeHelp.
@@
4. Results
4.1. RQ1: LLM Tutor Usage
@@
4.1.1. CodeHelp Enrollment and Usage
Table 1. CodeHelp Enrollment and Usage Rates for NNES and NES Students
@@ 4.1.1. CodeHelp Enrollment and Usage
Data on student enrollment and volume of queries asked can be found in Table 1.
Table 1 presents the enrollment rates for each group. To examine differences in student enrollment, we conducted a $\chi^{2}$ test comparing the overall class composition to the number of students enrolled with the CodeHelp system. The analysis yielded a $\chi^{2}$ statistic of 0.04 with a $p$-value greater than 0.05, indicating no significant difference in enrollment rates.
Table 1 also presents the usage rate (queries per student) for each group. To assess differences in CodeHelp use, a $\chi^{2}$ test was conducted by comparing the number of students enrolled in the CodeHelp system to the number of queries made. The analysis yielded a $\chi^{2}$ statistic of 16.89 with $p\ll 0.01$, indicating a significant difference in the number of queries made with NNES students making fewer queries. An effect size of $\varphi=0.13$ was calculated, suggesting a small effect.
4.1.2. Unique Users per Day per Hour
Figure 2. Unique Users per Day per Hour
@@ 4.1.2. Unique Users per Day per Hour
4.1.3. Language Use
Table 2. Non-English Languages Used in Queries
@@ 4.1.3. Language Use
Figure 3. Non-English Query Examples
The results of the languages used can be found in Table 2. From the 269 total queries from NNES students (Table 1), 25 of them were non-English, resulting in 9% of NNES queries being non-English. Additionally, we see NES students ask a few non-English queries as well. We note that many of the queries were multilingual in nature with English primarily being used for Java keywords. Figure 3 shows examples of the non-English queries that were submitted by students.
4.2. RQ2: LLM Tutor Views
4.2.1. General Views
@@
4.2.2. Help Seeking Preferences
Figure 4. Student Preferences for Help Resources
@@ 4.2.2. Help Seeking Preferences
4.2.3. LLM Tutor Evaluation
Table 3. Replicated Likert Questions from Liffiton et al. (Liffiton et al., 2023)
@@ 4.2.3. LLM Tutor Evaluation
We replicated the Likert questions posed to participants in the original CodeHelp study by Liffiton et al. (Liffiton et al., 2023) in our first survey. Our results can be found in Table 3. We found similar results with students typically finding CodeHelp useful to complete work and learn course material. Students were largely supportive of CodeHelp being available in future CS courses. To identify significant differences, we ran an independent samples t-test across all questions comparing NNES and NES responses. There were no significant differences for any question.
4.2.4. LLM Tutor Views
Table 4. Likert Questions in Survey 2 Compared Across NNES and NES
@@ 4.2.4. LLM Tutor Views
Using the open-ended responses from Survey 1, we created targeted questions for Survey 2. We ran independent samples t-tests on all questions, using the Šídák correction to adjust the significance level to account for multiple comparisons. All questions, along with their means, p-values, and corrected p-values, are listed in Table 4. Again, no significant differences were found between NNES and NES responses for any question (post-correction). In the initial independent samples t-tests, it was found that one question showed a significant difference between NNES and NES students at $p=0.02$ (“I find the quality of CodeHelp responses to be better than asking a tutor.”). Upon further inspection with the Šídák correction, we found that none of the comparisons were significantly different.
4.2.5. NNES Questions
Table 5. Likert Questions for NNES Students
@@ 4.2.5. NNES Questions
Specific questions for NNES were posed in an attempt to better understand their experience in computing and their LLM tutor usage. The questions and the results can be found in Table 5. We found that the NNES students in our study typically found it difficult to express and understand computing ideas in their native language, with both related questions having over 50% agreement. Their usage of CodeHelp depending on the language of the terminology is generally mixed. Lastly, we found that NNES students in our study typically do not know the programming technical terminology in their native language.
5. Discussion
5.1. NNES use CodeHelp less than NES
Our analysis revealed that NNES and NES students typically enrolled in the CodeHelp system at similar rates, which aligned with our expectations. However, contrary to our expectations, NNES students generally used CodeHelp less frequently than NES students (albeit with a small effect size). This is interesting coupled with the results found in help-seeking behaviors (Section 4.2.2) that showed NNES students have a similar preference for LLM tutors compared to NES students.
@@ 5.1. NNES use CodeHelp less than NES
We also found that NNES students used a larger variety of languages when interacting with CodeHelp compared to NES students. This result was expected, however it was notable to see NES students also using languages other than English. Additionally, many of the queries were multilingual using English computing keywords. This is in line with the result of NNES students not knowing computing terminology in their native language (Section 4.2.5).
5.2. NNES and NES evaluate CodeHelp similarly—but NNES rank online help higher
We saw that NNES students tended to rank online resources, such as CodeHelp, higher than NES students. This suggests that NNES students may be more proactive in seeking assistance online—possibly due to a greater need for supplementary resources to bridge language and comprehension gaps in traditional classroom settings. This communication apprehension and linguistic anxiety from NNES with others (such as friends and the instructor) is consistent with findings from past studies (Sobotka and Raman, 2020; Pappamihiel, 2001; Agarwal et al., 2022). Interestingly, we saw that both groups rated LLM tutors above LLMs. This is possibly due to having the LLM tutor “provided” by the course which telegraphs an endorsement by the professor for usage in the classroom (backed by finding in Section 4.2.1).
Both NNES and NES students evaluated the LLM tutor similarly, consistent with the results of the original CodeHelp study (Liffiton et al., 2023). This indicates that the LLM tutor, CodeHelp, was perceived as an effective educational tool across different studies. Additionally, students appreciated that CodeHelp had guardrails and guided to the answer—in line with previous findings on LLM tutor scaffolding (Denny et al., 2024b). By offering guidance instead of direct answers, it enables students to maintain autonomy in their learning. Lastly, we see that NNES students do not seem as comfortable with computing terminology in their native languages—in line with previous research reporting NNES student difficulties in translating this terminology (Alaofi and Russell, 2023). This discomfort may be attributed to the predominance of English in computer science education and industry (Hanselman, 2008; Chistyakov, 2017; Igawa et al., [n.d.]), resulting in limited exposure to technical terms in other languages.
6. Limitations

### v2

@@ I Introduction
In computing education, being a non-native English speaker (NNES) can present learning challenges, especially for those outside anglophone countries or with limited English proficiency, due to the predominance of English-based materials [5]. The most popular programming languages (e.g., Java, Python, C) use English for their keywords and API naming conventions [18, 9, 22]. English computer terminology used in classroom settings is one of the main factors negatively affecting NNES student performance in programming classes [26], predicting approximately 50% of final exam performance when controlling for English proficiency [2]. Moreover, English proficiency is strongly correlated with programming performance [35, 36, 41, 34], further widening the gap for NNES students. In addition to programming languages being primarily based in English, most online resources for learning programming (e.g., documentation, online Q&A forums, tutorials) are also predominantly in English. These skewed linguistic accommodations make a certain level of English proficiency increasingly necessary to learn computer programming [18]. This de facto prerequisite creates a significant hurdle for NNES individuals who may have limited English proficiency [17]. While tailoring the delivery of materials for NNES students has been suggested [6], the practical implementation of this strategy has traditionally been a challenge. However, large language models (LLMs) may now provide an opportunity to better serving NNES computing students as the models having been trained on 100+ natural languages [43]. Indeed, with the multilingual capabilities of LLMs and their past implementation in computing classrooms [15, 32, 28, 29, 42], this technology might help alleviate the de facto primarily English setting by allowing NNES students to easily obtain assistance in their native language.
@@
With the rapid rise of LLMs, computing educators face the challenge of integrating this technology into the classroom [7]. The study by Denny et al. suggests that LLMs can enhance the accessibility and scalability of high-quality educational content while reducing educators’ workloads [11]. One implementation of this technology is through tutoring systems that provide guidance, rather than direct answers to programming questions, helping users to reach solutions independently [25, 27].
@@
LLM tutoring systems have shown positive effects for students in the classroom [25, 27, 44]. Coupled with this, state-of-the-art LLMs that have been trained on 100+ languages [43] can perform extremely well when translating between languages [24]. Due to this, we hypothesize that NNES students may especially benefit from the use of LLM tutoring systems to ask for help—including in their native language. To understand this, we asked the following research questions:
  1. 1.
@@
II Related Work
II-A Non-Native English Speaker Stressors and LLM Assistance
@@
Past research has highlighted several challenges faced by NNES students in academic environments, particularly in fields like computer science [26, 18, 17, 3]. Alaofi and Russell highlighted various NNES student difficulties in CS1, including reading, understanding, or using error messages, understanding or using hints and tips from Integrated Development Environments in English, and general English language difficulties [3]. Due to a mixture of these difficulties, NNES students often experience linguistic anxiety [31] and communication apprehension [39] which can be exacerbated in an academic setting where technical terminology is prevalent. These hurdles, coupled with the embarrassment that NNES students often feel when asking for help [1], can impact their learning process and willingness to seek necessary assistance. Additionally, tutors have reported additional time required to support NNES students [21], which is in line with tutors reporting overall difficulty in implementing best tutor practices due to time constraints [37].
II-B LLMs in the Computing Classroom
@@
The increasing use of LLMs in computing classrooms [15, 32, 28, 29, 16, 42, 13] presents a valuable resource for students and new opportunities to support learning. State of the art models provide an alternative avenue for help seeking and can therefore aid learning while potentially reducing the workload on tutors [14, 29], although potential challenges to learning also exist [33]. Recent work also points to particular benefits for NNES students [4]. Within computing education, NNES students have been found to grasp computing concepts more quickly when using their native language [10]. Given the strong performance of LLMs in non-English languages [40, 23, 19] and the fact that many models are trained on over 100 languages [43], LLMs could be particularly beneficial for students studying in their preferred language. LLMs have the potential to effectively bridge language gaps in computing education, which is especially important given that learning to interact productively with LLMs is likely to become an essential skill for all students [12].
@@
This leads to our study in which we use the CodeHelp LLM tutor [27, 38]. CodeHelp is an open-source LLM-powered tool designed with guardrails to offer on-demand assistance to computing students without directly providing code solutions. We chose this tool as prior research indicated that it was well-received by students and instructors alike, and presented users with a simple interface [27]. CodeHelp was praised for its availability, help with resolving programming errors, easy deployment, and its supplemental support [27].
III Method
III-A LLM Tutor Deployment and Study Context
@@
In this study, we use the CodeHelp system—an open source LLM-powered tutor that includes built-in guardrails to produce general guidance toward a solution to a programming problem while being prevented from providing direct solutions [27]. Students interact with CodeHelp in a simple form, providing the programming language of interest (e.g., Java), an extract of the code related to the doubt, the error message, and the student’s question. The form can be seen in Figure 1. The written question is what will be analyzed in this study.
Figure 1: CodeHelp user interface illustrating input fields
@@
The implementation of CodeHelp aimed to enhance students’ learning experiences by providing an accessible, AI-driven resource for supplementary instruction and problem-solving support. CodeHelp leverages the OpenAI API with a choice to use the GPT-3.5-Turbo or the GPT-4o models. In our study, we used the GPT-3.5-Turbo model for financial reasons. CodeHelp was introduced to the course via a class forum post from the instructor. After some use with the system, we noticed that the CodeHelp system would always provide help in English, regardless of the language of the query. Due to the nature of this study, we coordinated with the CodeHelp developer to update the system to reply with the same natural language in which a question was asked, a characteristic known as language fidelity [20].
@@
Participants for this study were recruited from the pool of 321 students enrolled in the aforementioned accelerated course. All students in the course had access to the CodeHelp tool, but were not required to participate in our study surveys. CodeHelp was available to students for a total of five weeks. Each student independently signed up to the CodeHelp system, ensuring individualized tracking of interactions and usage patterns. Enrollment in the study was voluntary, and students were informed of the study’s purpose, procedures, and potential benefits prior to participating. In order to incentivize survey completion, the professor of the course offered extra credit points for completing surveys but students could still chose to participate in the study or not. In total, 80 students participated in the first survey and 162 students participated in the second survey, for a total of 170 unique students. Of these 170 students, 109 were native English speakers and 61 were non-native English speakers (as seen on Table I). All participation procedures were in accordance with our approved human subjects protocol.
III-B Survey Design and Administration
@@ III-C Data Analysis
To answer RQ1 around interactions with the LLM tutor, the enrollment and usage rates were analyzed using a $\chi^{2}$ test to explore differences between NNES and NES students. Specifically, we compared the proportions of NNES and NES students who enrolled in CodeHelp and those who actively used the platform post-enrollment. Additionally, we calculated the unique users by hour of a given day to analyze usage patterns for each group. For natural language use, the langdetectlangdetect Python package was used to identify the languages asked. A manual inspection of all non-English queries was also carried out to verify the correct language identification. We examined the frequency and provide descriptive statistics on the languages used to ask queries to CodeHelp.
@@
IV Results
IV-A RQ1: LLM Tutor Usage
@@
IV-A1 CodeHelp Enrollment and Usage
TABLE I: CodeHelp Enrollment and Usage Rates for NNES and NES Students
@@ IV-A1 CodeHelp Enrollment and Usage
Data on student enrollment and volume of queries asked can be found in Table I.
Table I presents the enrollment rates for each group. To examine differences in student enrollment, we conducted a $\chi^{2}$ test comparing the overall class composition to the number of students enrolled with the CodeHelp system. The analysis yielded a $\chi^{2}$ statistic of 0.04 with a $p$-value greater than 0.05, indicating no significant difference in enrollment rates.
Table I also presents the usage rate (queries per student) for each group. To assess differences in CodeHelp use, a $\chi^{2}$ test was conducted by comparing the number of students enrolled in the CodeHelp system to the number of queries made. The analysis yielded a $\chi^{2}$ statistic of 16.89 with $p\ll 0.01$, indicating a significant difference in the number of queries made with NNES students making fewer queries. An effect size of $\varphi=0.13$ was calculated, suggesting a small effect.
IV-A2 Unique Users per Day per Hour
Figure 2: Unique Users per Day per Hour
@@ IV-A2 Unique Users per Day per Hour
IV-A3 Language Use
TABLE II: Non-English Languages Used in Queries
@@ IV-A3 Language Use
Figure 3: Non-English Query Examples
The results of the languages used can be found in Table II. From the 269 total queries from NNES students (Table I), 25 of them were non-English, resulting in 9% of NNES queries being non-English. Additionally, we see NES students ask a few non-English queries as well. We note that many of the queries were multilingual in nature with English primarily being used for Java keywords. Figure 3 shows examples of the non-English queries that were submitted by students.
IV-B RQ2: LLM Tutor Views
IV-B1 General Views
@@
IV-B2 Help Seeking Preferences
Figure 4: Student Preferences for Help Resources
@@ IV-B2 Help Seeking Preferences
IV-B3 LLM Tutor Evaluation
TABLE III: Replicated Likert Questions from Liffiton et al. [27]
@@ IV-B3 LLM Tutor Evaluation
We replicated the Likert questions posed to participants in the original CodeHelp study by Liffiton et al. [27] in our first survey. Our results can be found in Table III. We found similar results with students typically finding CodeHelp useful to complete work and learn course material. Students were largely supportive of CodeHelp being available in future CS courses. To identify significant differences, we ran an independent samples t-test across all questions comparing NNES and NES responses. There were no significant differences for any question.
IV-B4 LLM Tutor Views
TABLE IV: Likert Questions in Survey 2 Compared Across NNES and NES
@@ IV-B4 LLM Tutor Views
Using the open-ended responses from Survey 1, we created targeted questions for Survey 2. We ran independent samples t-tests on all questions, using the Šídák correction to adjust the significance level to account for multiple comparisons. All questions, along with their means, p-values, and corrected p-values, are listed in Table IV. Again, no significant differences were found between NNES and NES responses for any question (post-correction). In the initial independent samples t-tests, it was found that one question showed a significant difference between NNES and NES students at $p=0.02$ (“I find the quality of CodeHelp responses to be better than asking a tutor.”). Upon further inspection with the Šídák correction, we found that none of the comparisons were significantly different.
IV-B5 NNES Questions
TABLE V: Likert Questions for NNES Students
@@ IV-B5 NNES Questions
Specific questions for NNES were posed in an attempt to better understand their experience in computing and their LLM tutor usage. The questions and the results can be found in Table V. We found that the NNES students in our study typically found it difficult to express and understand computing ideas in their native language, with both related questions having over 50% agreement. Their usage of CodeHelp depending on the language of the terminology is generally mixed. Lastly, we found that NNES students in our study typically do not know the programming technical terminology in their native language.
V Discussion
V-A NNES use CodeHelp less than NES
Our analysis revealed that NNES and NES students typically enrolled in the CodeHelp system at similar rates, which aligned with our expectations. However, contrary to our expectations, NNES students generally used CodeHelp less frequently than NES students (albeit with a small effect size). This is interesting coupled with the results found in help-seeking behaviors (Section IV-B2) that showed NNES students have a similar preference for LLM tutors compared to NES students.
@@ V-A NNES use CodeHelp less than NES
We also found that NNES students used a larger variety of languages when interacting with CodeHelp compared to NES students. This result was expected, however it was notable to see NES students also using languages other than English. Additionally, many of the queries were multilingual using English computing keywords. This is in line with the result of NNES students not knowing computing terminology in their native language (Section IV-B5).
V-B NNES and NES evaluate CodeHelp similarly—but NNES rank online help higher
We saw that NNES students tended to rank online resources, such as CodeHelp, higher than NES students. This suggests that NNES students may be more proactive in seeking assistance online—possibly due to a greater need for supplementary resources to bridge language and comprehension gaps in traditional classroom settings. This communication apprehension and linguistic anxiety from NNES with others (such as friends and the instructor) is consistent with findings from past studies [39, 31, 1]. Interestingly, we saw that both groups rated LLM tutors above LLMs. This is possibly due to having the LLM tutor “provided” by the course which telegraphs an endorsement by the professor for usage in the classroom (backed by finding in Section IV-B1).
Both NNES and NES students evaluated the LLM tutor similarly, consistent with the results of the original CodeHelp study [27]. This indicates that the LLM tutor, CodeHelp, was perceived as an effective educational tool across different studies. Additionally, students appreciated that CodeHelp had guardrails and guided to the answer—in line with previous findings on LLM tutor scaffolding [14]. By offering guidance instead of direct answers, it enables students to maintain autonomy in their learning. Lastly, we see that NNES students do not seem as comfortable with computing terminology in their native languages—in line with previous research reporting NNES student difficulties in translating this terminology [3]. This discomfort may be attributed to the predominance of English in computer science education and industry [18, 9, 22], resulting in limited exposure to technical terms in other languages.
VI Limitations